# BASS: Batched Attention-optimized Speculative Sampling

### v1

@@ 1 Introduction
To address this, we present Batched Attention-optimized Speculative Sampling (BASS) – a parallel speculative decoder that handles multiple sequences simultaneously. BASS increases GPU utilization by parallelism across both the batch dimension and the draft-token dimension. We implement customized CUDA kernels to handle ragged tensors during attention calculation, which are a challenge posed by batched speculative decoding, and design a heuristic to dynamically adjust draft length for each step. As illustrated in Figure 1, BASS achieves latency and GPU utilization that are substantially improved from prior regular and speculative decoders. By comprehensive experiments on three different models including CodeGen and OPT, we study these trends as well as accuracy benefits of BASS. We study the impact of draft model design on overall system performance as well as that of algorithmic choices in attention kernels and draft lengths.
@@ 2.1 Inference with LLMs
This paper focuses on transformer-based Vaswani et al. (2017) decoder-only generative LLMs. The standard inference of these LLMs can be divided into two phases: (a) Context encoding phase where the input prompt is processed in parallel to encode contextual information, and (b) Incremental decoding phase where the model auto-regressively generates output tokens one by one based on the encoded context (Figure 2(a)).
@@ 4.1 Setup
Models and tasks: We report experimental results on three main models with their respective draft models and tasks:
@@
Latency metrics: We use the metric of per-token latency for each generated sequence. For regular decoding, this value is the same across a batch. For speculative decoding, it varies within a batch and therefore we report three metrics: per-token latency of the first finished sequence in a batch, per-token latency of the last finished sequence in a batch, and per-token latency averaged across a batch, where each of the three is then averaged across examples in a task dataset.
4.2 OPT on XSum
Table 1 shows the accuracy and latency results of the OPT 13B model on the XSum task, with OPT 125M as the draft model. As expected, the results suggest neutral accuracy between regular decoding and speculative decoding, while speculative decoding provides up to 2.81$\times$ speed up for finishing the first sequence and up to 2.34$\times$ speed up on average for all sequences. In a real-life application, and particularly in the scenario of generating multiple sequences for the same prompt, we can respond to the user as soon as the first sequence finishes while the other additional recommendations continue to generate. Therefore, speeding up the first sequence by 2.05$\times$–2.81$\times$ implies a significant improvement in user-perceived latency.
@@ 4.5 Benefits of batched speculative decoding
Figure 5 emulates an application scenario where we have an end-to-end time budget of 2.5 seconds to generate 256-long sequences for any given prompt from HumanEval. According to Table 3, regular decoding under any setting would be unable to finish before time runs out, while single-sequence speculative decoding is able to return an answer with accuracy that corresponds to the left end point of the curves. As the batch size increases, the probability (Pass@Finished) that at least one returned sequence solves the coding task is increased up to 61%, and the probability (Pass@First) that we display a correct answer first to user is increased up to 43% with a simple ranking strategy using model confidence of mean-logP value. Both numbers are substantially higher than the mid-thirties accuracy by single-sequence speculative decoding. A real-life application would use a domain-specific stopping criteria instead of a fixed length and a more sophisticated ranking method, but the relative comparisons among the competing methods are as captured by Figure 5 and BASS is clearly superior.
@@ 4.6 Ablation studies
In Table 6, we compare latency when certain alternative implementation choices are used. With BASS-SPLIT, we launch per-sequence kernels to handle ragged tensors as illustrated in Figure 4(c). The results suggest that the cost of launching more CUDA kernels in BASS-SPLIT out-weights the cost of wasted compute in BASS-PAD. Note that this relation is task dependent and may change when the sequence lengths across a batch vary by a large margin. With “fixed draft size”, we use a constant draft length instead of Algorithm 1 that dynamically modifies draft length. The results suggest that both the efficient attention calculation and the draft-length heuristic are important to the performance of BASS.

### v2

@@ 1 Introduction
To address this, we present Batched Attention-optimized Speculative Sampling (BASS) – a parallel speculative decoder that handles multiple sequences simultaneously. BASS increases GPU utilization by parallelism across both the batch dimension and the draft-token dimension. We implement customized CUDA kernels to handle ragged tensors during attention calculation, which are a challenge posed by batched speculative decoding, and design a heuristic to dynamically adjust draft length for each step. As illustrated in Figure 1, BASS achieves latency and GPU utilization that are substantially improved from prior regular and speculative decoders. By comprehensive experiments on three different models including CodeGen and OPT, we study these trends as well as accuracy benefits of BASS. We study the impact of draft model design on overall system performance as well as that of algorithmic choices in attention kernels and draft lengths. BASS is applicable to both batch generation from a same prompt and batch generation from a set of different prompts.
@@ 2.1 Inference with LLMs
This paper focuses on transformer-based Vaswani et al. (2017) decoder-only generative LLMs. The standard inference of these LLMs can be divided into two phases: (a) context encoding (prefill) phase where the input prompt is processed in parallel to encode contextual information, and (b) incremental decoding phase where the model auto-regressively generates output tokens one by one based on the encoded context (Figure 2(a)).
@@ 3.2 Proposed method
The degree of alignment between draft and main models varies across prompts, across different sequences from the same prompt, and also within the same sequence. When generating commonly used sentences or boilerplate code for example, the alignment tends to be strong and the optimal draft length is long. When generating uncommon sentences or novel code segments, the alignment tends to be weak and the optimal draft length is short. The effect of Algorithm 1 versus a fixed draft length is to dynamically get closer to the optimal draft length in both scenarios, so that the system generates longer drafts where possible yet does not waste compute to generate throw-away tokens.
@@ 4.1 Setup
Models and tasks: We report experimental results on three main models with their respective draft models and tasks [^5] While the three tasks are the scenario of batch generation from a same prompt, please note that BASS is also applicable to batch generation from a set of different prompts.:
@@
Latency metrics: We use the metric of per-token latency for each generated sequence  [^6] It is important to note that we do not divide latency by batch size, which was done in some papers, e.g., Su et al. (2023). Fundamentally, our definition is a latency metric while the definition in Su et al. (2023) is a throughput metric. With our definition, per token latency increases as the batch size increases because the amount of FLOPS during the per token latency is multiplied by batch size, and this applies to both regular decoding and speculative decoding.. For regular decoding, this value is the same across a batch. For speculative decoding, it varies within a batch and therefore we report three metrics: per-token latency of the first finished sequence in a batch, per-token latency of the last finished sequence in a batch, and per-token latency averaged across a batch, where each of the three is then averaged across examples in a task dataset.
4.2 Performance on summarization
Table 1 shows the accuracy and latency results of the OPT 13B model on the summarization task of the XSum dataset, with OPT 125M as the draft model. As expected, the results suggest neutral accuracy between regular decoding and speculative decoding, while speculative decoding provides up to 2.81$\times$ speed up for finishing the first sequence and up to 2.34$\times$ speed up on average for all sequences. In a real-life application, and particularly in the scenario of generating multiple sequences for the same prompt, we can respond to the user as soon as the first sequence finishes while the other additional recommendations continue to generate. Therefore, speeding up the first sequence by 2.05$\times$–2.81$\times$ implies a significant improvement in user-perceived latency.
@@ 4.5 Benefits of batched speculative decoding
Figure 5 emulates a real-life application scenario where a service returns code recommendations to a user within a time budget. One of the recommendations is first displayed and the user has the option to flip through others. Ranking (here simply mean-logP based) is applied to pick the first displayed one. The Pass@First metric is the probability that the first displayed recommendation solves the problem correctly. The Pass@Finished metric is the probability that at least one of the finished recommendations within the time budget solves the problem correctly. These two metrics together quantify the accuracy quality of the service.
Figure 5 uses an end-to-end time budget of 2.5 seconds to generate 256-long sequences for any given prompt from HumanEval. According to Table 3, regular decoding under any setting would be unable to finish before time runs out, while single-sequence speculative decoding is able to return one recommendation and its Pass@First and Pass@Finished are the same and correspond to the left end point of the curves. With BASS and as the batch size increases, Pass@Finished is increased up to 61% and Pass@First is increased up to 43% with a simple ranking strategy using model confidence of mean-logP value. Both numbers are substantially higher than the mid-thirties accuracy by single-sequence speculative decoding. A real-life application would use a domain-specific stopping criteria instead of a fixed length and a more sophisticated ranking method, but the relative comparisons among the competing methods are as captured by Figure 5 and BASS is clearly superior.
@@ 4.6 Ablation studies
In Table 6, we compare latency when certain alternative implementation choices are used. With BASS-SPLIT, we launch per-sequence kernels to handle ragged tensors as illustrated in Figure 4(c). The results suggest that the cost of launching more CUDA kernels in BASS-SPLIT out-weights the cost of wasted compute in BASS-PAD. Note that this relation is task dependent and may change when the sequence lengths across a batch vary by a large margin. For example, when applied on tasks of batch generation from a set of different prompts, the advantages of BASS-SPLIT could out-weights the cost. With “fixed draft size”, we use a constant draft length instead of Algorithm 1 that dynamically modifies draft length. The results suggest that both the efficient attention calculation and the draft-length heuristic are important to the performance of BASS.
@@
Code 7.8B, HumanEval 1st Seq PTL (ms)
batch 2 4 8
BASS 3.7 4.1 4.5
BASS-SPLIT 4.0 4.4 5.2
fixed draft size 4 4.6 5.1 6.3
fixed draft size 6 4.3 4.9 5.8
fixed draft size 8 4.0 4.3 5.1